# EifBench: Extremely Complex Instruction Following Benchmark for Large Language Models

### v1

@@ Abstract
Tao Zou, Xinghua Zhang, Haiyang Yu, Minzheng Wang, Fei Huang, Yongbin Li thanks: *Corresponding author Tongyi Lab, Alibaba Group ✉ : {qingdie.zt, zhangxinghua.zxh, yifei.yhy, wangminzheng.wmz, f.huang, shuide.lyb}@alibaba-inc.com
@@ 1 Introduction
The advent of large language models (LLMs) has transformed real-world applications by improving models’ ability to comprehend a diverse range of human instructions, from simple conversations to complex problem solving Sanh et al. (2022); Dubois et al. (2023). Thus, instructions have become central to effective human-machine interaction in this new landscape Zhong et al. (2021); Mishra et al. (2022); Gao et al. (2024), especially the paradigm of “Model as Product” has deeply entered the collective consciousness where LLM agents need to accurately complete a series of tasks to meet user demands Xiong et al. (2025); Hu et al. (2024); Alakuijala et al. (2025). However, as user demands grow more sophisticated, traditional benchmarks Zhong et al. (2024); Chia et al. (2023), which focus on specific tasks, are insufficient to evaluate models’ comprehensive ability to handle multifaceted instructions. This shortfall underscores the need for innovative evaluation frameworks capable of accurately assessing how models understand and execute complex instructions Zhou et al. (2023); Wang et al. (2023); Xu et al. (2024).
@@
  • We conduct a detailed analysis of 18 LLMs, encompassing both open-source and closed-source models, uncovering their limitations in processing complex instructions and pinpointing areas for enhancement to better adapt to real-world complex scenarios. The SegPO algorithm demonstrates significant improvements, achieving increases of 14.85% compared to the base LLM and 3.40% compared to GRPO models on EifBench, respectively.
@@ 2.2 Construction Workflow
2) Multi-scenario Data Collection. Our collection process involves three types of datasets: plain text, dyadic dialogue, and multi-party dialogue. Plain text samples are drawn from existing works Wen et al. (2024); Li et al. (2024). For dyadic dialogues, we gather real-life interactions, undergo cleaning and noise reduction, and use large language models (LLMs) to condense conversations while preserving key information. Multi-party dialogue data is synthesized with LLMs, crafting diverse scenarios and participant numbers. Specific prompts guide LLMs to produce varied and representative dialogue content, enhancing the depth and applicability of the dataset.
@@ 2.3 Evaluation Protocol
We employ GPT-4o OpenAI (2023) as the evaluation model to assess constraint adherence in generated responses. Following established practices Wen et al. (2024), the $k$-th constraint in the $j$-th instruction of $i$-th instance is given a binary compliance score $S_{i,j,k}\in\{0,1\}$, with 1 signifying full compliance and 0 indicating non-compliance.
@@
$\text{ILA}_{i}$ $=\frac{1}{m_{i}}\sum_{j=1}^{m_{i}}\left(\frac{1}{c_{i,j}}{\prod_{k=1}^{c_{i,j}}S_{i,j,k}}\right)$ (1)
@@ 3 Preliminaries
Existing Large Reasoning Models (LRMs) Reid et al. (2024); Jaech et al. (2024) have demonstrated improved performance on complex tasks via structured reasoning. A pivotal element in training these LRMs is the use of the Group Relative Policy Optimization (GRPO) Shao et al. (2024) algorithm in Reinforcement Learning (RL). Given a query $q$, GRPO samples a group of outputs $\{o_{1},o_{2},\cdots,o_{G}\}$ from the old policy $\pi_{{\theta}_{old}}$ and eliminates the need for a separate value function and instead relies on the average reward as a baseline to compute the advantage. The optimization process for policy is as follows:
@@ 5.1 Baselines
We compare the performance of both proprietary and open-source LLMs trained on diverse corpora. In the proprietary category, we evaluate models such as GPT-4o OpenAI (2023), GPT-4o-mini OpenAI (2023), Claude3.5-Sonnet Anthropic (2024b), Claude3.5-Haiku Anthropic (2024a), gemini-1.5-Pro Reid et al. (2024) and GPT-o3-mini. Among open-source models, we assess LLaMA3.1 Dubey et al. (2024), Qwen2 Yang et al. (2024a), Qwen2.5, DeepSeek-R1 Reid et al. (2024), QwQ-32B Yang et al. (2024b), and Qwen3 Yang et al. (2025) to explore their efficiency.
@@ 5.2 Settings
For inference, we efficiently process proprietary models through their APIs. For open-source models, we employ a robust setup consisting of four Nvidia A100 GPUs, each equipped with 80GB of VRAM, utilizing the vLLM framework on EifBench where applicable. This configuration enables the completion of all tasks in roughly 30 minutes. During evaluation, the GPT-4o model serves as the evaluator, with assessment durations ranging from 4 to 10 hours based on task complexity. Our code and dataset are available at https://github.com/Hope-Rita/EIFBench.
@@ 5.3.1 How do existing LLMs perform?
Model performance varies notably across categories, revealing task-type dependencies. In closed-source models, GPT-4o excels in ILA with relatively lower CLA. This indicates its capability to focus and complete individual sub-tasks effectively, albeit less so on fulfilling comprehensive constraints. In the realm of open-source models, Qwen2.5-72B-Instruct performs exceptionally well, balancing both instruction completion and constraint adherence without requiring much thought. Among reasoning models, DeepSeek-R1 and Qwen3-32B demonstrate robust competitiveness, effectively handling complex reasoning tasks. These findings emphasize the varying strengths of models across different task types and their alignment with specific task demands.
@@ C.1 Factors Influencing Instruction Following
C.2 Performance from state-of-art LLMs
We employ other state-of-the-art large language models, i.e., gemini-2.0-Flash, as external evaluators to assess performance. The results are shown in Table 9. Their evaluations generally align with those of GPT-4o-as-judge, demonstrating consistent outcomes.
Model Plain Text Dyadic Dialogue Multi-party Dialogues
Closed-Source LLMs
GPT-4o 0.3248 0.6941 0.2273 0.6100 0.2641 0.6155
Claude-3.5-Sonnet 0.1283 0.4235 0.0800 0.4860 0.0988 0.4200
GPT-4o-mini 0.1263 0.5898 0.0892 0.5588 0.1197 0.6432
Claude-3.5-Haiku 0.0463 0.2425 0.0365 0.2379 0.0295 0.1288
gemini-1.5-Pro 0.2001 0.7018 0.1732 0.7349 0.1920 0.7747
GPT-o3-mini 0.4443 0.8706 0.3112 0.7503 0.4302 0.9002
Open-Source LLMs
LLaMA3.1-8B-Instruct 0.0293 0.3954 0.0184 0.3357 0.0149 0.4151
LLaMA3.1-70B-Instruct 0.0502 0.4616 0.0439 0.5337 0.0451 0.5086
Qwen2-7B-Instruct 0.0475 0.4371 0.0346 0.4237 0.0321 0.4731
Qwen2-72B-Instruct 0.1300 0.6688 0.1272 0.6847 0.1173 0.6941
Qwen2.5-7B-Instruct 0.0767 0.5677 0.0654 0.6224 0.0570 0.6310
Qwen2.5-72B-Instruct 0.2182 0.7656 0.1732 0.7395 0.1678 0.7887
QwQ-32B-Preview 0.1017 0.5203 0.0640 0.4632 0.0840 0.5659
DeepSeek-R1 0.2449 0.6985 0.1797 0.7390 0.2060 0.7412
Qwen3-32B 0.2826 0.8002 0.2702 0.7924 0.2894 0.8328
Qwen3-32B w/o think 0.2724 0.8035 0.2726 0.7929 0.2559 0.8246
Qwen3-235B-A22B 0.2382 0.6949 0.2624 0.7527 0.2499 0.7518
Qwen3-235B-A22B w/o think 0.2435 0.6951 0.2571 0.7442 0.2645 0.7651
Table 9: Performance metrics across different task categories: Plain Text, Dyadic Dialogue, and Multi-party Dialogue with gemini-2.0-Flash as evaluator. The best and second-best results are highlighted in bold and underlined.

### v2

@@ Abstract
Tao Zou, Xinghua Zhang, Haiyang Yu, Minzheng Wang, Fei Huang, Yongbin Li thanks: *Corresponding author Tongyi Lab, Alibaba Group {qingdie.zt, zhangxinghua.zxh, yifei.yhy, wangminzheng.wmz, f.huang, shuide.lyb}@alibaba-inc.com
@@ 1 Introduction
The advent of large language models (LLMs) has transformed real-world applications by improving models’ ability to comprehend a diverse range of human instructions, from simple conversations to complex problem solving Sanh et al. (2022); Dubois et al. (2023); Zhang et al. (2024c). Thus, instructions have become central to effective human-machine interaction in this new landscape Zhong et al. (2021); Mishra et al. (2022); Gao et al. (2024), especially the paradigm of “Model as Product” has deeply entered the collective consciousness where LLM agents need to accurately complete a series of tasks to meet user demands Xiong et al. (2025); Hu et al. (2024); Alakuijala et al. (2025). However, as user demands grow more sophisticated, traditional benchmarks Zhong et al. (2024); Chia et al. (2023), which focus on specific tasks, are insufficient to evaluate models’ comprehensive ability to handle multifaceted instructions. This shortfall underscores the need for innovative evaluation frameworks capable of accurately assessing how models understand and execute complex instructions Zhou et al. (2023); Wang et al. (2023); Xu et al. (2024).
@@
  • We conduct a detailed analysis of 20 LLMs, encompassing both open-source and closed-source models, uncovering their limitations in processing complex instructions and pinpointing areas for enhancement to better adapt to real-world complex scenarios. The SegPO algorithm demonstrates significant improvements, achieving increases of 14.85% compared to the base LLM and 3.40% compared to GRPO models on EifBench, respectively.
@@ 2.2 Construction Workflow
2) Multi-scenario Data Collection. Our collection process involves three types of datasets: plain text, dyadic dialogue, and multi-party dialogue. Plain text samples are drawn from existing works Wen et al. (2024); Li et al. (2024). For dyadic dialogues, we first perform data cleaning and noise reduction on the collected real-life interactions. Guided by the methodology in Wang et al. (2025b), we employ Large Language Models (LLMs) to condense the conversations while ensuring the preservation of key information. Multi-party dialogue data is synthesized with LLMs, crafting diverse scenarios and participant numbers. Specific prompts guide LLMs to produce varied and representative dialogue content, enhancing the depth and applicability of the dataset.
@@ 2.3 Evaluation Protocol
We employ GPT-4o OpenAI (2023) as the evaluation model to assess constraint adherence in generated responses in LLM-as-judge manner Zhang et al. (2023); Wang et al. (2024b); Zeng et al. (2024); Wang et al. (2024a). Following established practices Wen et al. (2024), the $k$-th constraint in the $j$-th instruction of $i$-th instance is given a binary compliance score $S_{i,j,k}\in\{0,1\}$, with 1 signifying full compliance and 0 indicating non-compliance.
@@
$\text{ILA}_{i}$ $=\frac{1}{m_{i}}\sum_{j=1}^{m_{i}}{\prod_{k=1}^{c_{i,j}}S_{i,j,k}}$ (1)
@@ 3 Preliminaries
Existing Large Reasoning Models (LRMs) Wang et al. (2025a); Jaech et al. (2024) have demonstrated improved performance on complex tasks via structured reasoning. A pivotal element in training these LRMs is the use of the Group Relative Policy Optimization (GRPO) Shao et al. (2024) algorithm in Reinforcement Learning (RL). Given a query $q$, GRPO samples a group of outputs $\{o_{1},o_{2},\cdots,o_{G}\}$ from the old policy $\pi_{{\theta}_{old}}$ and eliminates the need for a separate value function and instead relies on the average reward as a baseline to compute the advantage. The optimization process for policy is as follows:
@@ 5.1 Baselines
We compare the performance of both proprietary and open-source LLMs trained on diverse corpora. In the proprietary category, we evaluate models such as GPT-4o OpenAI (2023), GPT-4o-mini OpenAI (2023), Claude3.5-Sonnet Anthropic (2024b), Claude3.5-Haiku Anthropic (2024a), gemini-1.5-Pro Reid et al. (2024), gemini-2.0-Flash and GPT-o3-mini. Among open-source models, we assess LLaMA3.1 Dubey et al. (2024), Qwen2 Yang et al. (2024a), Qwen2.5, DeepSeek-V3 DeepSeek-AI et al. (2024), DeepSeek-R1 Reid et al. (2024), QwQ-32B Yang et al. (2024b), and Qwen3 Yang et al. (2025) to explore their efficiency.
@@ 5.2 Settings
For inference, we efficiently process proprietary models through their APIs. For open-source models, we employ a robust setup consisting of four Nvidia A100 GPUs, each equipped with 80GB of VRAM, utilizing the vLLM framework on EifBench where applicable. This configuration enables the completion of all tasks in roughly 30 minutes. During evaluation, the GPT-4o model serves as the evaluator, with assessment durations ranging from 4 to 10 hours based on task complexity. Our code and dataset are publicly available for reproducibility. [^3] Available at https://github.com/Hope-Rita/EIFBench and https://github.com/Tongyi-CCAI/Complex-IF.
@@ 5.3.1 How do existing LLMs perform?
Model performance varies notably across categories, revealing task-type dependencies. In closed-source models, GPT-4o excels in ILA with relatively lower CLA. This indicates its capability to focus and complete individual sub-tasks effectively, albeit less so on fulfilling comprehensive constraints. Well-performing models in the open-weight landscape generally fall into two categories. The first is generalist models, like Qwen2.5-72B-Instruct, where reasoning emerges from large-scale and instruction tuning. Specialized models like DeepSeek-R1 and Qwen3-32B, equipped with explicit reasoning strategies, demonstrate their advanced capabilities with a distinct performance profile. They achieve substantially superior performance on ILA metrics, while also showing robust, comparable performance on CLA metrics against their non-reasoning counterparts. This highlights that their specialized architecture enables a deeper understanding of complex tasks, leading to success on targeted metrics.
@@ 7 Conclusion
Acknowledgments
We would like to thank the reviewers for their helpful reviews and feedback. This work was supported by Alibaba Research Intern Program.